# Towards deterministic non-Gaussianity on a chip

### v1

@@ Abstract
We propose cascaded spontaneous four-wave mixing (SFWM) in microring resonators as a scalable and efficient approach for directly generating non-Gaussian states of light. Focusing on the well-understood “low-gain” regime, we demonstrate that triplet generation through cascaded SFWM can be achieved with high efficiency and favorable spectral characteristics using realistic microring sources in AlGaAs. The ability to achieve the generation of light in a single set of supermodes – and the accessibility of the “high-gain” regime at realistic pump powers – makes this source a promising candidate as a direct and deterministic source of non-Gaussian light for photonic quantum information processing.
@@
Introduction—Generating non-Gaussian states of light is a central topic in the quantum optics and photonic quantum information processing communities. Much of the work has been motivated by underlying fundamental physics: Due to their uniquely “non-classical” characteristics, non-Gaussian states are a necessary resource for implementing photonic quantum information processing [1]. But creating non-Gaussian states is challenging [2]. The conventional approach is to apply measurements – usually photon-number-resolving (PNR) detection – to squeezed light [3, 4, 5]. Although PNR detectors and squeezed light sources are readily available, relying on detection to introduce non-Gaussianity has the drawbacks of being probabilistic, introducing latency times, and requiring cryogenic components in a setting that could otherwise operate at room temperature. This has led to a search for deterministic, room-temperature sources of non-Gaussian light based on the use of parametric processes [6].
@@
Another proposal relies on pump depletion and the resulting entanglement of pump light with photon pairs generated by the pump. This has been theoretically considered in the so-called mesoscopic regime in scenarios involving long waveguides [9, 10], and in microring resonators requiring both very high finesse and pump powers [11]. Experimental implementation of these schemes have been hampered by fabrication requirements and material limitations, and issues of scalability and few-supermode generation can also be of concern.
@@
An alternative strategy would be the use of cascaded squeezing, where a squeezed vacuum is generated and then used to drive another squeezed light source; because the second process is driven by a non-classical pump, it has the desired non-Gaussian form [9, 6]. If the first squeezing interaction is restricted to the low-gain regime, this results in the probabilistic generation of photon triplets: This process – and the certification of the resulting tripartite entanglement – has been implemented through spontaneous parametric down-conversion (SPDC) in bulk crystals [12, 13, 14, 15] and in integrated waveguides [16]. Besides the inherent scalability, the use of integrated structures offers an enhanced effective nonlinearity due to the strong transverse field confinement.
@@
Yet implementations of cascaded squeezing in microring resonators have not yet been investigated in detail, despite the fact that microring sources satisfy the three requirements mentioned above. First, they provide an additional resonant field enhancement, further increasing the structure’s effectiveness in light generation [17]. We will also show that the type of “few-supermode” operation we seek can be obtained very effectively using resonant sources, often achieving a higher level of separability than non-resonant sources [7, 18]. Finally, the scalability of microring resonators has already made them a popular choice in the implementation of large-scale integrated devices.
@@
The enhanced nonlinearity in these devices allows one to envision implementing cascaded spontaneous four-wave mixing (SFWM) rather than SPDC. This has a number of practical advantages: The energy conservation conditions in SFWM allow for all the fields to be very close to each other in frequency, for instance, in the same telecom band. Phase matching – which can be a significant challenge in triplet generation – is therefore widely accessible. The narrow frequency range should also make it relatively easy to find appropriate material platforms that are transparent at the relevant wavelengths and that do not exhibit deleterious effects such as two-photon absorption.
@@
For the sample calculations, we envision the use of integrated AlGaAs structures, but the scheme we propose can be implemented in other platforms. As well, to benchmark the results that can be expected in first experiments, we focus on the probabilistic generation of photon triplets, although with reasonable experimental parameters one can go beyond that limit.
@@
Cascaded spontaneous four-wave mixing—We propose a source consisting of two rings (see Fig. 1a), such that the first (second) pair generation process occurs in the first (second) ring: This is ensured by tuning the positions of the ring resonances, resulting in a flexible and practical design in which each ring’s resonance frequencies, free spectral range (FSR), and waveguide coupling can be controlled independently. Although single-ring cascaded SFWM could be implemented in principle, this would be challenging, and it would require careful dispersion engineering to ensure the proper resonance alignment. A device with two rings also allows for the two pump fields to be confined to separate rings, preventing some of the parasitic processes that would occur if both interactions occurred in a single ring. The basic microring structures sketched in Fig. 1a are sufficient for demonstrations of cascaded SFWM, but in principle they could be replaced by more sophisticated structures – e.g., with interferometric couplers or photonic molecules – to provide even more tunability of individual resonances, and suppression of unwanted nonlinear processes [19, 20, 21, 22].
@@
We envision working at telecom wavelengths, and we focus on a scheme in which the generated modes are non-degenerate in frequency (see Fig. 2a). We choose this configuration to clearly identify the different modes involved in the interactions, but the design can be modified to affect the positions of the pump and generated fields; for example, degenerate triplet sources can be designed by adjusting the positions of the rings’ resonances (see Fig. 2b).
@@
Figure 1: a) Subsequent microring resonators coupled to a bus waveguide, each coupled to a phantom waveguide to model photon losses. b) Electric field ($E_{x}$) mode profile of the TE00 mode. Mode simulations are done in Lumerical with a bend radius of 10 $\mu\mathrm{m}$, using the refractive index for Al0.3Ga0.7As from [23].
@@
The Hamiltonian describing the nonlinearity is given by [24]
$H_{\mathrm{NL}}=-\frac{1}{4\epsilon_{0}}\int\mathrm{d}\bm{r}\Gamma^{(3)}_{ijkl}(\bm{r})D^{i}(\bm{r})D^{j}(\bm{r})D^{k}(\bm{r})D^{l}(\bm{r}),$ (1)
where repeated Cartesian indices are implicitly summed over and where $\bm{D}(\bm{r})$ denotes the displacement field in the structure [24], which we expand in terms of asymptotic fields [25, 26] having the form
$\bm{D}_{u}^{\mathrm{in}}(\bm{r})=\sum_{\lambda}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\mathrm{in},\lambda}(\bm{r})a_{u}^{\mathrm{in},\lambda}(k)+\mathrm{H.c.,}$ (2)
and similarly for the “out” fields, and we show the linear Hamiltonian in Appendix Towards deterministic non-Gaussianity on a chip. The index $\lambda$ denotes an input or output channel through which light can propagate [27, 28], and for the structure given in Fig. 1a, $\lambda\in\{\mathrm{ac,ph(1),ph(2)}\}$ for the actual, first, and second phantom channels; the phantom channels are used to account for scattering losses [25]. The $a_{u}^{\text{in/out},\lambda}(k)$ denote bosonic lowering operators with the usual commutation relations [27], and $\bm{\mathcal{D}}_{uk}^{\text{in/out},\lambda}(\bm{r})$ denotes the asymptotic field amplitude. The form of the latter depends on the structure in question (see Supplemental Material 1).
We restrict ourselves to the terms in the full nonlinear Hamiltonian relevant to cascaded SFWM: $H_{\mathrm{NL}}^{(I)}(t)=H_{\mathrm{SFWM1}}^{(I)}(t)+H_{\mathrm{SFWM2}}^{(I)}(t)$ (see Appendix Towards deterministic non-Gaussianity on a chip). The state generated by the nonlinear interactions is given by
@@
The final term in Eq. (4) refers to the generation of a photon triplet through cascaded SFWM. The spectral correlations for a photon triplet are characterized by the triphoton wavefunction (TWF), given in Eq. (32).
@@
Few-supermode operation—The ability to generate light in the minimal number of supermodes is essential in developing a useful source of non-Gaussianity, just as one seeks single-Schmidt-mode sources of squeezed light [29, 7, 30, 31, 32]. In the low-gain regime this is achieved by obtaining a TWF that is separable [7], and although more work is needed to understand non-Gaussian nonlinear optics in the high gain regime, this approach can be expected to be useful there as well.
@@
Sample calculation—We now compute triplet generation TWFs and rates for a realistic device. We consider implementations in the AlGaAs platform: This emerging platform supports large nonlinear coefficients $\gamma_{\mathrm{NL}}$, has demonstrated high quality factors [34], and is compatible with hybrid integration approaches [35]. In our calculations, we take $\chi^{(3)}\approx 5.6\times 10^{-19}$ m2/V2 ($n_{2}\approx 1.43\times 10^{-17}$ m2/W) [36, 34, 35], corresponding to nonlinear parameters of $\gamma_{\mathrm{NL}}\approx 250\,(\mathrm{W\cdot m})^{-1}$ for both SFWM processes, computed using the TE00 modes (see Fig. 1b and Supplemental Material 2). For the ring dimensions we take $R_{1}={20\,\mu\mathrm{m}}$ and $R_{2}={10\,\mu\mathrm{m}}$ (recall the labelling in Fig. 1a), both with cross-sectional dimensions of $(400\times 920)$ nm (see Fig. 1b). We assume realistic intrinsic quality factors of $Q_{\mathrm{int}}=10^{6}$ for all the resonances in the frequency range of interest [34]. The rings and pumps should be tuned such that both pumps are on resonance.
We consider the pump fields $P_{1}$ and $P_{2}$ for the first and second ring respectively (recall Fig. 1 & 2), to be centered at wavelengths $\lambda_{P_{1}}=1546.83$ nm and $\lambda_{P_{2}}=1584.29$} nm, respectively. The triplets are generated in the resonances with center wavelengths $\lambda_{S_{1}}=1541.60$ nm, $\lambda_{S_{2}}=1562.68$ nm, and $\lambda_{S_{3}}=1573.41$ nm. The rate of triplets at the chip’s output is given by $R_{3}=R_{R}\eta_{S_{1}}^{\mathrm{ac}(1)}\eta_{S_{2}}^{\mathrm{ac}(2)}\eta_{S_{3}}^{\mathrm{ac}(2)}|\sigma|^{2}$, where $R_{R}$ is the pulses’ repetition rate, and $\eta_{u}^{\lambda(i)}$ are the escape efficiencies (see Eq. (17)).
@@
Initially we consider two configurations $A$ and $B$, with critical coupling for all resonances, such that all loaded quality factors are $5\times 10^{5}$. We take $P_{1}$ to be a Gaussian pulse with full-width-half-maximum (FWHM) of $300$ ps – similar to the ring’s dwelling time – with a $R_{R}=10$ MHz repetition rate. Interesting triplet rates can be obtained with either pulsed or CW excitation in the second ring. In configuration $A$, $P_{2}$ is CW; in configuration $B$, it is a train of $300$ ps pulses. To obtain $|\beta|^{2}\approx 0.1$, we limit $P_{1}$ to a peak power of $0.53\,\mathrm{mW}$ in the bus waveguide. In configuration $A$, we take the power of the CW $P_{2}$ to be 1.4{1.4} mW, which corresponds to $\sim 1$ W circulating in the ring; this results in an output triplet rate of $R_{3}\approx{7.43}$ Hz. In configuration $B$, the pulsed $P_{2}$ is set such that the peak power in the ring does not exceed 10 W; this results in a peak power in the waveguide of 0.12 W, giving $R_{3}=19.2$ Hz.
We represent the TWFs obtained with configurations $A$ and $B$ in Fig. 3 (a) and (b), respectively. In moving from configuration $A$ to configuration $B$, we see that – similar to photon-pair generation in a ring resonator [40] – pumping with a pulse duration shorter than the dwelling time in the ring results in a TWF with reduced spectral correlations. The visualizations of the TWFs in Fig. 3 are isosurfaces at 10% of the modulus squared of the maximal value of the given TWF, and projections over each variable are provided on the panels.
@@
Discussion—Our proposed device demonstrates that photon triplet generation can be achieved on a platform that is simultaneously efficient, compact, and readily manufacturable using existing fabrication techniques. Our proposed design requires lower average pump powers compared to other proposals, requiring only average on-chip powers of $\sim$ 1 mW or less, and we have used well-accepted material parameters. Furthermore, our resonant photon-triplet source offers both spectral tunability and control over the number of supermodes, providing essential flexibility for tailoring the properties of the generated non-Gaussian states. Our proposed device could obtain triplet rates comparable to those quoted in other theoretical proposals [42, 43, 44, 45], while being more efficient, compact, and allowing for tunability of the light’s spectral properties. For comparison to experimentally demonstrated sources, taking a loss of 3 dB per output channel [46, 47] for a total loss of 9 dB; this would result in a measured triplet coincidence rate of $\sim 1$ Hz, higher than the rates observed in experimental implementations of cascaded pair generation [12, 13, 16, 14]. In addition, we note that the proposed device could achieve higher triplet rates by further increasing the repetition rate, since the resonator’s dwelling times (maximally $\sim 400$ ps for critical coupling) do not limit the repetition rate to the 10 MHz used in our calculations. We believe these high rates and promising characteristics strongly motivate the fabrication and demonstration of these sources, including the characterization of the generated light, which could be done through stimulated emission tomography [48], for example.
Perhaps most interesting is the fact that this source, due to its high efficiency, is a strong candidate for implementing cascaded SFWM in a high-gain regime: Above, we restricted the first SFWM process to a low-gain regime by restricting the energy of $P_{1}$ to less than a picojoule. From an experimental perspective, one can easily envision increasing the strength of $P_{1}$, resulting in squeezed vacuum which would then drive the second SFWM process, with the possibility of measurable non-Gaussianity in the field generated from the latter process. The dynamics of cascaded SFWM in this high-gain regime, and the characteristics of the deterministically generated non-Gaussian state it would produce, will be a rich and challenging topic for future theoretical work. It will also be a novel source of non-classical light in more complex quantum photonic devices [6, 49, 50].

### v2

# Photon triplets from integrated microrings: A path towards deterministic non-Gaussianity on a chip

@@ Abstract
We propose cascaded spontaneous four-wave mixing (SFWM) in microring resonators as a scalable and efficient approach for directly generating non-Gaussian states of light. Focusing on the well-understood “low-gain” regime, we demonstrate that triplet generation through cascaded SFWM can be achieved with high efficiency and favorable spectral characteristics using realistic microring sources in AlGaAs. The ability to achieve the generation of light in a single set of supermodes – and the predicted accessibility of the “high-gain” regime at realistic pump powers – makes this source a promising candidate as a direct and deterministic source of non-Gaussian light for photonic quantum information processing.
@@
I Introduction
Generating non-Gaussian states of light is a central topic in the quantum optics and photonic quantum information processing communities. Due to their uniquely “non-classical” characteristics, non-Gaussian states are a necessary resource for implementing photonic quantum information processing [42]. But creating non-Gaussian states is challenging [51]. The conventional approach is to apply measurements – usually photon-number-resolving (PNR) detection – to squeezed light [44, 45, 28]. Although PNR detectors and squeezed light sources are readily available, relying on detection to introduce non-Gaussianity has the drawbacks of being probabilistic, introducing latency times, and requiring cryogenic components in a setting that could otherwise operate at room temperature. This has led to a search for deterministic, room-temperature sources of non-Gaussian light based on the use of parametric processes [53].
@@ I Introduction
Another proposal relies on pump depletion and the resulting entanglement of pump light with photon pairs generated by the pump. This has been theoretically considered in the so-called mesoscopic regime in scenarios involving long waveguides [54, 52], and in microring resonators requiring very high finesse and very high pump powers [47]. Experimental implementation of these schemes has been hampered by fabrication requirements and material limitations, and issues of scalability and few-supermode generation can also be of concern.
@@
An alternative strategy would be the use of cascaded squeezing, where a squeezed vacuum is generated and then used to drive another squeezed light source; because the second process is driven by a non-classical pump, it has the desired non-Gaussian form [54, 53]. If the first squeezing interaction is restricted to the low-gain regime, this results in the probabilistic generation of photon triplets: This process – and the certification of the resulting tripartite entanglement – has been implemented through spontaneous parametric down-conversion (SPDC) in bulk crystals [21, 16, 29, 11] and in integrated waveguides [27], and in cascaded quantum dot schemes [25]. Besides the inherent scalability, the use of fully integrated structures offers an enhanced effective nonlinearity due to the strong transverse field confinement.
@@
Yet implementations of cascaded squeezing in microring resonators have not yet been investigated in detail, despite the fact that microring sources satisfy the three requirements mentioned above. First, they provide an additional resonant field enhancement, further increasing the structure’s effectiveness in light generation and, importantly, increasing the brightness by having comparable or higher generation rates within a smaller bandwidth than non-resonant integrated sources [19]. We will also show that the type of “few-supermode” operation we seek can be obtained very effectively using resonant sources, often achieving a higher level of separability than non-resonant sources [32, 20]. Finally, the scalability of microring resonators has already made them a popular choice in the implementation of large-scale integrated devices.
@@
The enhanced nonlinearity in these devices allows one to envision implementing cascaded spontaneous four-wave mixing (SFWM) rather than SPDC. This has a number of practical advantages: The energy conservation conditions in SFWM allow for all the fields to be very close to each other in frequency, for instance in the same telecom band. Phase matching – which can be a significant challenge in triplet generation – is therefore widely accessible. The narrow frequency range should also make it relatively easy to find appropriate material platforms that are transparent at the relevant wavelengths and that do not exhibit deleterious effects such as two-photon absorption.
@@
For the sample calculations, we envision the use of integrated AlGaAs structures, but the scheme we propose can be implemented in other platforms. As well, to benchmark the results that can be expected in first experiments, we focus on the probabilistic generation of photon triplets. We will also show that one can go beyond that limit with reasonable experimental parameters.
II Cascaded spontaneous four-wave mixing
@@
We propose a source consisting of two rings (see Fig. 1a), such that the first (second) pair generation process occurs in the first (second) ring: This is ensured by tuning the positions of the ring resonances, resulting in a flexible and practical design in which each ring’s resonance frequencies, free spectral range (FSR), and waveguide coupling can be controlled independently. Although single-ring cascaded SFWM could be implemented in principle, this would be challenging, and it would require careful dispersion engineering to ensure the proper resonance alignment. A device with two rings also allows for the two pump fields to be confined to separate rings, preventing some of the parasitic processes that would occur if both interactions occurred in a single ring. The basic microring structures sketched in Fig. 1a are sufficient for demonstrations of cascaded SFWM, but in principle they could be replaced by more sophisticated structures – e.g., with interferometric couplers or photonic molecules – to provide even more tunability of individual resonances, and suppression of unwanted nonlinear processes [49, 34, 55, 40].
@@
We envision working at telecom wavelengths, and we focus on a scheme in which the generated modes are non-degenerate in frequency (see Fig. 2a). We choose this configuration to clearly identify the different modes involved in the interactions, but the design can be modified to affect the positions of the pump and generated fields; for example, degenerate triplet sources can be designed by adjusting the positions of the ring resonances (see Fig. 2b).
@@
Figure 1: a) Subsequent microring resonators coupled to a bus waveguide, each coupled to a phantom waveguide to model photon losses. The $\hat{z}=\hat{z}^{\prime}$ unit vectors are perpendicular to the plane of the chip. b) Electric field ($E_{x}$) mode profile of the TE00 mode. Mode simulations are done in Lumerical with a bend radius of 10 $\mu\mathrm{m}$, using the refractive index for Al0.3Ga0.7As from Adachi et al. [1].
@@
The Hamiltonian describing the nonlinearity is ([39])
$H_{\mathrm{NL}}=-\frac{1}{4\epsilon_{0}}\int\mathrm{d}\boldsymbol{r}\Gamma^{(3)}_{ijkl}(\boldsymbol{r})D^{i}(\boldsymbol{r})D^{j}(\boldsymbol{r})D^{k}(\boldsymbol{r})D^{l}(\boldsymbol{r}),$ (1)
where repeated Cartesian indices are implicitly summed over and where $\boldsymbol{D}(\boldsymbol{r})$ denotes the displacement field in the structure [39], which we expand in terms of asymptotic fields [30, 7] having the form
$\boldsymbol{D}_{u}^{\mathrm{in}}(\boldsymbol{r})=\sum_{\lambda}\int\mathrm{d}k\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{in},\lambda}(\boldsymbol{r})a_{u}^{\mathrm{in},\lambda}(k)+\mathrm{H.c.,}$ (2)
and similarly for the “out” fields. We show the explicit expressions for the linear and nonlinear Hamiltonians in Appendices A and B, respectively. The tensor $\Gamma^{(3)}$ is related to the usual third-order nonlinear susceptibility tensor $\chi^{(3)}$ [39], as presented in Appendix C (see Eq. (40)). In Eq. (2), the index $\lambda$ denotes an input or output channel through which light can propagate [5, 15], and for the structure given in Fig. 1a, $\lambda\in\{\mathrm{ac,ph(1),ph(2)}\}$ for the actual, first, and second phantom channels; the phantom channels are used to account for scattering losses [30]. The $a_{u}^{\text{in/out},\lambda}(k)$ denote bosonic lowering operators with the usual commutation relations [5], and $\boldsymbol{\mathcal{D}}_{uk}^{\text{in/out},\lambda}(\boldsymbol{r})$ denotes the asymptotic field amplitude. The form of the latter depends on the structure in question, and we detail those for the cascaded rings from Fig. 1a in Appendix D.
We restrict ourselves to the terms in the full nonlinear Hamiltonian relevant to cascaded SFWM: $H_{\mathrm{NL}}^{(I)}(t)=H_{\mathrm{SFWM1}}^{(I)}(t)+H_{\mathrm{SFWM2}}^{(I)}(t)$ (see Appendix B). The state generated by the nonlinear interactions is given by
@@
The final term in Eq. (4) refers to the generation of a photon triplet through cascaded SFWM. The spectral correlations for a photon triplet are characterized by the triphoton wavefunction (TWF), given in Eq. (32). We note that this triplet term can only arise due to the non-Gaussian nature of the cascaded SFWM Hamilitonian (Eq. (25)); therefore, a demonstration of triplet generation can be considered a signature of the state’s non-Gaussianity.
@@
III Few-supermode operation
The ability to generate light in the minimal number of supermodes is essential in developing a useful source of non-Gaussianity, just as one seeks single-Schmidt-mode sources of squeezed light [57, 32, 26, 17, 8]. In the low-gain regime this is achieved by obtaining a TWF that is separable [32], and although more work is needed to understand non-Gaussian nonlinear optics in the high gain regime, this approach can be expected to be useful there as well.
@@ III Few-supermode operation
IV Sample calculation
We now compute triplet generation TWFs and rates for a realistic device. We consider implementations with AlGaAs structures: This emerging platform supports large nonlinear coefficients $\gamma_{\mathrm{NL}}$, has demonstrated high quality factors [43], and is compatible with hybrid integration approaches [3]. In our calculations, we take $\chi^{(3)}\approx 5.6\times 10^{-19}$ m2/V2 ($n_{2}\approx 1.43\times 10^{-17}$ m2/W) [2, 43, 3], corresponding to nonlinear parameters of $\gamma_{\mathrm{NL}}\approx 250\,(\mathrm{W\cdot m})^{-1}$ for both SFWM processes, computed using the TE00 modes (see Fig. 1b and Appendix C). For the ring dimensions we take $R_{1}={20\,\mu\mathrm{m}}$ and $R_{2}={10\,\mu\mathrm{m}}$ (recall the labelling in Fig. 1a), both with cross-sectional dimensions of $(400\times 920)$ nm (see Fig. 1b). We assume realistic intrinsic quality factors of $Q_{\mathrm{int}}=10^{6}$ for all the resonances in the frequency range of interest [43]. The rings and pumps should be tuned such that both pumps are on resonance.
We consider the pump fields $P_{1}$ and $P_{2}$ for the first and second ring respectively (recall Fig. 1 & 2), to be centered at wavelengths $\lambda_{P_{1}}=1546.83$ nm and $\lambda_{P_{2}}=1584.29$} nm, respectively. The triplets are generated in the resonances with center wavelengths $\lambda_{S_{1}}=1541.60$ nm, $\lambda_{S_{2}}=1562.68$ nm, and $\lambda_{S_{3}}=1573.41$ nm. The rate of triplets at the chip’s output is given by $R_{3}=R_{R}\eta_{S_{1}}^{\mathrm{ac}(1)}\eta_{S_{2}}^{\mathrm{ac}(2)}\eta_{S_{3}}^{\mathrm{ac}(2)}|\sigma|^{2}$, where $R_{R}$ is the pulse repetition rate, and $\eta_{u}^{\lambda(i)}$ are the escape efficiencies (see Eq. (17)).
@@ IV Sample calculation
Initially we consider two configurations $A$ and $B$, with critical coupling for all resonances, such that all loaded quality factors are $5\times 10^{5}$. We take $P_{1}$ to be a Gaussian pulse with full-width-half-maximum (FWHM) of $300$ ps – similar to the ring dwelling time – with a repetition rate $R_{R}=10$ MHz. Interesting triplet rates can be obtained with either pulsed or CW excitation in the second ring. In configuration $A$, $P_{2}$ is CW; in configuration $B$, it is a train of $300$ ps pulses. To obtain $|\beta|^{2}\approx 0.1$, we limit $P_{1}$ to a peak power of $0.53\,\mathrm{mW}$ in the bus waveguide. In configuration $A$, we take the power of the CW $P_{2}$ to be 1.4{1.4} mW, which corresponds to $\sim 1$ W circulating in the ring; this results in an output triplet rate of $R_{3}\approx{7.43}$ Hz. In configuration $B$, the pulsed $P_{2}$ is set such that the peak power in the ring does not exceed 10 W; this results in a peak power in the waveguide of 0.12 W, giving $R_{3}=19.2$ Hz.
We represent the TWFs obtained with configurations $A$ and $B$ in Fig. 3a and 3b, respectively. In moving from configuration $A$ to configuration $B$, we see – similar to photon-pair generation in a ring resonator [18] – that pumping with a pulse duration shorter than the dwelling time in the ring results in a TWF with reduced spectral correlations. The visualizations of the TWFs in Fig. 3 are isosurfaces at 10% of the modulus squared of the maximal value of the given TWF, and projections over each variable are provided on the panels. The projections on the panels of Fig. 3 can be thought of as the spectral correlations between the two photons over which the TWF has not been integrated. For example, comparing the top right panels from Fig. 3a-c, we see a change in the spectral correlations between the photons in $S_{2}$ and $S_{3}$ (the perpendicular axis being the $S_{1}$ variable, over which the integration is performed), going from more correlated to less correlated, which can be seen by the relative thickness of the off-diagonal (similar to what is seen in photon pair generation [14]). Similarly, the overall shape of the three-dimensional isosurfaces can describe the correlations between the three photons constituting the triplet.
@@
We note that higher generation rates, brightnesses and purities can be achieved by tuning various pumping and coupling parameters. In this manuscript, our focus is on showing that high generation rates are possible for a range of purities. We leave the determination of the maximal generation rates and optimal purities to future work.
@@
V Discussion
Our proposed device demonstrates that photon triplet generation can be achieved on a platform that is simultaneously efficient, compact, and readily manufacturable using existing fabrication techniques. Our proposed design requires lower average pump powers compared to other proposals, requiring only average on-chip powers of $\sim$ 1 mW or less, and we have used well-accepted material parameters. Furthermore, it offers both spectral tunability and control over the number of supermodes, providing essential flexibility for tailoring the properties of the generated non-Gaussian states. It would yield triplet rates comparable to those quoted in other theoretical proposals [10, 35, 48, 4], while being more efficient, compact, and allowing for tunability of the light’s spectral properties. For comparison to experimentally demonstrated sources, taking a loss of 3 dB per output channel [9, 38] for a total loss of 9 dB; this would result in a measured triplet coincidence rate of $\sim 1$ Hz, higher than or similar to the rates observed in experimental implementations of cascaded pair generation [21, 16, 27, 29] (<< 1 Hz) and cascaded quantum dots [25] ($\sim$1 Hz). In addition, we note that the proposed device could achieve higher triplet rates by further increasing the repetition rate, since the resonator dwelling times (maximally $\sim 400$ ps for critical coupling) do not limit the repetition rate to the 10 MHz used in our calculations. We believe these high rates and promising characteristics strongly motivate the fabrication and demonstration of these sources, including the characterization of the generated non-Gaussian light, which could be done through stimulated emission tomography [13], for example. In the low-gain regime considered here, genuine three-photon coincidences would constitute sufficient evidence that the emitted light possesses non-Gaussian features.
Perhaps most interesting is the fact that this source, due to its high efficiency, is a strong candidate for implementing cascaded SFWM in a high-gain regime: Here, we restricted the first SFWM process to a low-gain regime by restricting the energy of $P_{1}$ to less than a picojoule. From an experimental perspective, one can easily envision increasing the strength of $P_{1}$, resulting in squeezed vacuum which would then drive the second SFWM process, where in this high-gain regime measurements of quadrature non-Gaussian features, such as Wigner negativity, in phase space [51] can be envisioned. The dynamics of cascaded SFWM in this high-gain regime, and the characteristics of the deterministically generated non-Gaussian state it would produce, will be a rich and challenging topic for future theoretical work. It will also be a novel source of non-classical light in more complex quantum photonic devices [53, 56, 12].